OpenStudy (anonymous):

Solve the equation: h(x)= -x^2 +8x-16

OpenStudy (anonymous):

х=4, maybe))

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

h'(x)= -2x+8 -2x+8=0 -2x=-8 x=4

OpenStudy (anonymous):

maybe

OpenStudy (merchandize):